The following is the information from the national income accounts for ahypothetical country:GDPRs. 6000.00Gross Investment 800.00Net Investment200.00Consumption 4000.00Govt. purchases of goods & services 1100.00Govt. Budget Surplus 30.00What is a) NDP b) Net exports c) Govt. taxes minus transfersc) Disposable personal income e) Personal Saving
1
Expert's answer
2016-11-14T09:59:12-0500
GDP = 6000, GI = 800, NI = 200, C = 4000, G = 1100, BS = 30,
a) NDP = GDP - NI + GI = 6600.
b) Net exports NE = GDP - C - G - GI = 6000 - 4000 - 1100 - 800 = 100.
c) Govt. taxes minus transfers T = G + BS = 1100 + 30 = 1130.
d) Disposable personal income DI = GDP - T = 6000 - 1130 = 4870.
e) Personal Saving S = GDP - C = 6000 - 4000 = 2000.